# Juan A. Rodriguez1,2,4 Abhay Puri1 Shubham Agarwal1, 2 Issam H. Laradji1, 5 Pau Rodriguez6* Sai Rajeswar1,2 David Vazquez1 Christopher Pal1,2,3 Marco Pedersoli4

### v3

@@ Abstract
1ServiceNow Research 2Mila - Quebec AI Institute 3Canada CIFAR AI Chair 4ÉTS, Montréal, Canada
5UBC, Vancouver, Canada 6Apple MLR, Barcelona, Spain * External collaboration
@@ 1 Introduction
Figure 2: (Left) Image Vectorization results using StarVector-8B, LIVE, and VTracer. Each row shows the input image, generated SVGs, and pixel-wise difference maps to highlight accuracy. StarVector-8B better preserves shapes, color gradients, and text, despite minor misplacements. Notably, MSE often misaligns with visual quality, e.g., regarding the ‘planet’ example, StarVector’s MSE (0.009) is higher than LIVE’s (0.0012) and VTracer’s (0.0039), yet StarVector preserves the color gradient. For the ‘diagram’ example, StarVector preserves the text. DinoScore better reflects these details, consistently favoring StarVector. (Right) Curve vs Primitive-based Vectorization. SVG code generated by StarVector and VTracer for the given image. StarVector effectively leverages shape primitives, resulting in a compact and precise vectorization. VTracer decomposes the image into numerous paths, resulting in a more complex result with less semantic clarity.
@@
Previous vectorization approaches fall into two main categories: traditional image processing methods and deep learning (DL)-based. Image processing methods [87, 51, 59] employ pixel-level analysis to trace vector curves, but often produce overly complex representations with artifacts and lack semantic understanding (Figure 2). While DL approaches [14, 67, 13, 89] have advanced vector graphics modeling through latent variable models and differentiable rendering [43, 50], they typically struggle with generalization beyond specific domains and underutilize SVG primitives. This limits their effectiveness for complex SVGs like scientific diagrams and precludes their use in modern multimodal tasks such as text-driven SVG generation [65, 72, 96, 24].
@@
We introduce [Uncaptioned image] StarVector, a foundational MLLM for SVG generation. StarVector processes both images and text instructions to produce compilable SVG code, leveraging SVG primitives to accurately represent vector graphics. We build upon the StarCoder works [42, 49] to connect the code generation research with SVG generation. Figure 3 describes the model architecture. It integrates an image encoder that projects images into visual tokens, and a transformer language model for learning the relationships between instructions, visual features, and SVG code sequences, to perform image vectorization (Image-to-SVG) or text-driven SVG Generation (Text-to-SVG) tasks. StarVector, performs primitive-aware vectorization through learned semantic understanding, effectively leveraging SVG primitives without explicit pixel reconstruction objectives. To address the lack of large-scale SVG datasets for training StarVector, we introduce [Uncaptioned image] SVG-Stack, containing over 2M SVG samples paired with rendered images and textual descriptions.
Additionally, we find that conventional metrics like MSE fail to adequately assess vector graphics fidelity, as demonstrated in Figure 2. Instead, we propose DinoScore, a perceptual similarity metric that better correlates with human perception of visual quality, and introduce [Uncaptioned image] SVG-Bench, a comprehensive evaluation framework spanning 10 datasets and 3 tasks: Image-to-SVG, Text-to-SVG, and diagram generation.
@@ 2 Related Work
SVG Generation Methods. Early efforts on image vectorization [92] predominantly utilized traditional image processing techniques [44, 95, 21, 44] such as segmentation and polynomial curve fitting [43, 87, 59, 51]. With the rapid advances in deep learning, new approaches have emerged. For example, SVG-VAE [47], a class-conditional Variational Autoencoder (VAE) [34], predicts a latent style vector and generates SVGs using an LSTM decoder [30]. DeepSVG [14] proposes a hierarchical VAE architecture using transformers to represent SVG paths. Im2Vec [67] translates pixel images into latent representations, which can be decoded into paths via a recurrent neural network. These methods are limited to generating basic primitives like paths; hence, they exhibit a performance gap with respect to image processing methods. Despite this, they open the door to more complex tasks that require understanding the semantic features of SVGs.
Recent trends in image generation using diffusion [29, 73] or autoregressive [23, 64, 96] models and the success of text-to-image generation [65, 72, 6] have inspired the research into text-conditioned SVG generation. VectorFusion [33] leverages a strong text-to-image diffusion model to find the SVG via iterative optimization. Some work has focused on SVG editing from textual inputs [12]. CLIPasso [86] uses a CLIP distance loss to iteratively refine SVG from sketches. These solutions are slow due to their iterative nature. IconShop [94] trains a BERT [20] model for text-conditioned SVG generation of icons, but their method is restricted to using paths.
@@
SVG Benchmarks and Datasets. Previous work on SVG datasets and benchmarks has been limited. Existing Image-to-SVG datasets mainly focus on fonts, icons, and emojis [14, 67, 13, 94, 12, 89], offering limited variety for broader SVG types and primitives. For text-driven SVG Generation, recent efforts [16, 94, 12] use large language models [56] to generate synthetic captions, but these datasets are mostly inaccessible, hindering reproducibility. Evaluation methods also face challenges. Existing pixel-based metrics like MSE [14, 67, 47] fail to capture the fidelity and structure of SVGs, overlooking aspects like line definition and primitive usage (Figure 2, illustrates how MSE can be misleading). Recent benchmarking efforts [54, 99] focus on caption-based generation or editing but remain limited in scope.
@@ 4.1 Architecture
    StarVector-1B is initialized with a CLIP ViT-B/32 [62] image encoder that processes images at a 224×224 resolution, producing 257 visual tokens. The LLM uses StarCoder-1B [42], with a context length of 8192 tokens.
@@
    StarVector-8B employs a SigLip (siglip-so400m-patch14-384) [97] image encoder that processes images at 384×384 resolution, yielding 576 visual tokens. This model utilizes StarCoder2-7B, which offers an expanded context length of 16k tokens. With a total of 8B parameters, we explore how scaling can produce more precise and compact SVGs due to the higher image resolution and enhanced LLM capacity.
@@
In response to the limited benchmarks available for SVG evaluation, and to unify evaluation practices [47, 14, 67], we introduce SVG-Bench. This benchmark evaluates SVG methods, focused on Image-to-SVG, Text-to-SVG, and Diagram Generation tasks. It covers a range of SVG types, from simple graphics like icons and fonts to complex structures like diagrams with numerous primitives. Dataset statistics are shown in Table 5.
@@ 5.1 Tasks and Benchmarks
    Image-to-SVG: This task evaluates converting images into SVGs. To account for varying visual complexities, we define the datasets SVG-Fonts, SVG-Emoji, SVG-Icons, and SVG-Stack, which provide increasing levels of complexity. Previous works have used fonts [47], emojis [67], and icons [14], but often with unclear access. We provide train, validation, and test splits for these datasets. We also include simplified versions of these datasets to support methods that only process simplified SVGs (containing only paths).
@@
Evaluation Metrics To compute benchmark scores, we define the following metrics: For image vectorization tasks, we use Mean Squared Error (MSE), Structural Similarity Index (SSIM) [91, 90], and Learned Perceptual Image Patch Similarity (LPIPS) [98]. To address the limitations of pixel-based metrics (see Figure 2), we propose DinoScore [57], which computes L2 distance between DinoV2 features. Token Length (Tokens) measures the size of the SVG samples. We use the StarCoder [42] tokenizer to tokenize SVG code and compute the average length. These metrics are also used for Diagram Generation. For Text-to-SVG, we build on text-to-image literature [81, 62, 66, 73] and prior Text-to-SVG methods [94, 12], using FID [81], FID-CLIP [94], and CLIP Score [62] to measure image-text alignment.
@@ 6 Experiments and Results
Table 2: Image Vectorization on SVG-Diagrams. StarVector leads in DinoScore, LPIPS, and SSIM, while LIVE excels in MSE. However, visual results (Fig. 2) show StarVector as the only effective method, highlighting MSE misalignment. It is also competitive in token length (Tokens).
@@
Baselines. In our comparisons, we consider the following model baselines. On Image-to-SVG, we compare the best image processing algorithms available, like Potrace [59], Vtracer [87], and Autotrace [51]. We report the performance of DL methods for Image-to-SVG, namely DeepSVG [14], Im2Vec [67], and MLLMs like GPT4 Vision [56]. For Text-to-SVG we consider IconShop, DeepSVG+GAN, and Bert, as presented in [94]. We also evaluate LLMs like CodeLlama-70b [82] and GPT-4 [55]. More details are available in Appendix 9, and Table 6 summarizes the SVG capabilities of all methods.
Training and Inference. We train StarVector-1B with a batch size of 128 and StarVector-8B with 512, using a learning rate of 1e-5 and AdamW optimizer. StarVector-1B took 7 days on 8 A100 GPUs, and StarVector-8B took 10 days on 64 H100 GPUs, both completing 2 epochs. Full training details are in Appendix 11.2. During inference, we generate $k=5$ SVG outputs with varying temperatures from 0 to 1 and select the sample with the highest DinoScore. We set a logit bias of 10 for the <svg-end> token, and apply top-p nucleus sampling (0.9) with a length penalty of -0.5. We use beam search with a beam size of 1 and employ vLLM [37] as the generation backend to significantly accelerate the generation process.
@@ 6.1 Main Results
Diagram Generation. Table 2 shows results on SVG-Diagrams, and Figures 4 and 14 provide visual examples. Results highlight that StarVector is the only method capable of performing diagram generation, as it uniquely applies the required primitives like rectangles, arrows, and text, whereas other methods produce blobs and curves that attempt to replicate structure and color. Mtrics such as DinoScore, LPIPS, capture this advantage, while MSE and SSIM remain poorly aligned. Human evaluations further confirmed the preference for StarVector’s outputs.
@@ 7 Conclusions
Limitations and Future Work. StarVector is limited by its 16k token context, which is insufficient for highly complex SVGs. It primarily relies on code prediction with minimal visual feedback, yet achieves high-quality results. Generation speed is also constrained by the LLM. Future work could explore integrating pixel-based signals to address these limitations.
@@ 8.3 Creating SVG-Diagrams
Figure 6: SVG Diagrams examples. These are ground truth SVG examples from the test set. They are presented as SVG, showing the challenge in understanding intricate structures and small texts, with images of variate aspect ratios.

### v4

# Juan A. Rodriguez1,2,4 Abhay Puri1 Shubham Agarwal1, 2 Issam H. Laradji1, 5 Pau Rodriguez1 Sai Rajeswar1,2 David Vazquez1 Christopher Pal1,2,3 Marco Pedersoli1,4

@@ Abstract
1ServiceNow Research 2Mila - Quebec AI Institute 3Canada CIFAR AI Chair
4ÉTS, Montréal, Canada 5UBC, Vancouver, Canada
@@ 1 Introduction
Figure 2: (Left) Image Vectorization results using StarVector-8B, LIVE, and VTracer. Each row shows the input image, generated SVGs, and pixel-wise difference maps to highlight accuracy. StarVector-8B better preserves shapes, color gradients, and text, despite minor misplacements. Notably, MSE often misaligns with visual quality, e.g., regarding the ‘planet’ example, StarVector’s MSE (0.009) is higher than LIVE’s (0.0012) and VTracer’s (0.0039), yet StarVector preserves the color gradient. For the ‘diagram’ example, StarVector preserves the text. DinoScore better reflects these details, consistently favoring StarVector. (Right) Curve vs Primitive-based Vectorization. SVG code generated by StarVector and VTracer for the given image. StarVector effectively leverages shape primitives, resulting in a compact vectorization. VTracer decomposes the image into numerous paths, resulting in a more complex result with less semantic clarity.
@@
Vectorization approaches fall into two main categories: traditional image processing and deep learning (DL)-based methods. Image processing techniques [87, 51, 59] trace vector curves via pixel-level analysis but often generate overly complex representations with artifacts and lack semantic understanding (Figure 2). DL methods [14, 67, 13, 89] improve vector modeling using latent variable models and differentiable rendering [43, 50] but struggle with generalization and underutilize SVG primitives. This limits their effectiveness for complex SVGs like scientific diagrams and hinders their application in modern multimodal tasks such as text-driven SVG generation [65, 72, 96, 24].
@@
We introduce [Uncaptioned image] StarVector, a foundational MLLM for SVG generation. StarVector processes both images and text instructions to produce compilable SVG code, leveraging SVG primitives to accurately represent vector graphics. We build upon the StarCoder works [42, 49] to connect the code generation research with SVG generation. Figure 3 describes the model architecture. It integrates an image encoder that projects images into visual tokens, and a transformer language model for learning the relationships between instructions, visual features, and SVG code sequences, to perform image vectorization (Image-to-SVG) or text-driven SVG Generation (Text-to-SVG) tasks. StarVector performs primitive-aware vectorization through learned semantic understanding, effectively leveraging SVG primitives without explicit pixel reconstruction objectives. To address the lack of large-scale SVG datasets for training StarVector, we introduce SVG-Stack, with over 2M SVG samples paired with rendered images and text descriptions.
Additionally, we find that conventional metrics like MSE fail to adequately assess vector graphics quality, as demonstrated in Figure 2. Instead, we propose DinoScore, a perceptual similarity metric that better correlates with visual human perception, and introduce SVG-Bench, an evaluation framework spanning 10 datasets and 3 tasks: Image-to-SVG, Text-to-SVG, and diagram generation.
@@ 2 Related Work
SVG Generation Methods. Early image vectorization efforts [92] primarily relied on traditional techniques [44, 95, 21] such as segmentation and polynomial curve fitting [43, 87, 59, 51]. With advancements in deep learning, new approaches have emerged. For instance, SVG-VAE [47] employs a class-conditional Variational Autoencoder (VAE) [34] to predict a latent style vector and generate SVGs using an LSTM decoder [30]. DeepSVG [14] introduces a hierarchical VAE architecture using transformers for SVG path representation, while Im2Vec [67] converts pixel images into latent representations that can be decoded into paths with a recurrent neural network. However, these methods are limited to basic primitives like paths, resulting in a performance gap compared to traditional image processing methods. Concurrently, BeyondPixels [99] utilizes LLMs to predict command sequences in synthetic SVGs. These trends highlight the potential of LLM models for improved generalization in SVG generation.
Recent results in image generation using diffusion [29, 73] or autoregressive [23, 64, 96] models and the success of text-to-image generation [65, 72, 6] have inspired the research into text-conditioned SVG generation. VectorFusion [33] leverages a strong text-to-image diffusion model to find the SVG via iterative optimization. Some work has focused on SVG editing from textual inputs [12]. CLIPasso [86] uses a CLIP distance loss to iteratively refine SVG from sketches. These solutions are slow due to their iterative nature. IconShop [94] trains a BERT [20] model for text-conditioned SVG generation of icons, but their method is restricted to using paths.
@@
SVG Benchmarks and Datasets. Previous work on SVG datasets and benchmarks has been limited. Existing Image-to-SVG datasets mainly focus on fonts, icons, and emojis [14, 67, 13, 94, 12, 89], offering limited variety for broader SVG types and primitives. Recent text-driven SVG generation efforts [16, 94, 12] leverage large language models [56] for synthetic captions, but limited dataset accessibility hampers reproducibility. Evaluation methods also face challenges. Existing pixel-based metrics like MSE [14, 67, 47] fail to capture the fidelity and structure of SVGs, overlooking aspects like line definition and primitive usage (Figure 2, illustrates how MSE can be misleading). Recent benchmarking efforts [54, 100] focus on caption-based generation or editing but remain limited in scope.
@@ 4.1 Architecture
    StarVector-1B is initialized with a CLIP ViT-B/32 [62] image encoder, processing images at a 224×224 resolution to produce 257 visual tokens. The LLM uses StarCoder-1B [42] with a context length of 8192 tokens.
@@
    StarVector-8B employs a SigLip (siglip-so400m-patch14-384) [97] image encoder, processing images at a 384×384 resolution to yield 576 visual tokens. This model utilizes StarCoder2-7B, offering an expanded context length of 16k tokens. With a total of 8B parameters, we investigate how scaling can yield more precise and compact SVGs due to higher image resolution and enhanced LLM capacity.
@@
In response to the limited benchmarks for SVG evaluation and to unify evaluation practices [47, 14, 67], we introduce SVG-Bench. This benchmark assesses SVG methods across Image-to-SVG, Text-to-SVG, and Diagram Generation tasks, encompassing various SVG types from simple graphics like icons and fonts to complex diagrams with multiple primitives. Dataset statistics are presented in Table 5.
@@ 5.1 Tasks and Benchmarks
    Image-to-SVG: This task evaluates converting images to SVGs across varying complexities. We introduce SVG-Fonts, SVG-Emoji, SVG-Icons, and SVG-Stack, increasing in complexity. While prior works used similar datasets [47, 67, 14], access has often been unclear. We provide standardized train, validation, and test splits, along with simplified versions containing only paths for compatibility with certain methods.
@@
Evaluation Metrics.
To compute benchmark scores, we define the following metrics: For image vectorization tasks, we use Mean Squared Error (MSE), Structural Similarity Index (SSIM) [91, 90], and Learned Perceptual Image Patch Similarity (LPIPS) [98]. To address the limitations of pixel-based metrics (see Figure 2), we propose DinoScore [57], which computes L2 distance between DinoV2 features. Token Length (Tokens) measures the size of the SVG samples. We use the StarCoder [42] tokenizer to tokenize SVG code and compute the average length. These metrics are also used for Diagram Generation. For Text-to-SVG, we build on text-to-image literature [81, 62, 66, 73] and prior Text-to-SVG methods [94, 12], using FID [81], FID-CLIP [94], and CLIP Score [62] to measure image-text alignment.
@@ 6 Experiments and Results
Table 2: Image Vectorization on SVG-Diagrams. StarVector outperforms LIVE in DinoScore, LPIPS, and SSIM, while LIVE ranks best in MSE. However, visual results (Fig. 2) confirm that StarVector is the only effective method for SVG generation, underscoring the misalignment of MSE. Additionally, it remains competitive in terms of token length.
@@
Baselines. Baselines. For our comparisons, we consider the following model baselines: For Image-to-SVG, we evaluate top image processing algorithms such as Potrace [59], Vtracer [87], and Autotrace [51]. We also report on deep learning methods including DeepSVG (<5M parameters) [14], Im2Vec (<5M) [67], and MLLMs like GPT-4 Vision (>100B) [56]. For Text-to-SVG, we consider methods like IconShop (>1B), DeepSVG+GAN (<5M), and BERT (>1B), as outlined in [94]. Additionally, we evaluate LLMs such as CodeLlama-70b [82] and GPT-4 (>100B) [55]. For more details, see Appendix 9, and Table 6 summarizes the SVG capabilities of all methods.
Training and Inference. We train StarVector-1B with a batch size of 128 and StarVector-8B with 512, using a learning rate of 1e-5 and the AdamW optimizer. StarVector-1B took 7 days on 8 A100 GPUs, while StarVector-8B took 10 days on 64 H100 GPUs, both completing 2 epochs. Full training details are in Appendix 11.2. During inference, we generate $k=5$ SVG outputs with temperatures ranging from 0 to 1, selecting the one with the highest DinoScore. We set a logit bias of 10 for the <svg-end> token and apply top-p nucleus sampling (0.9) with a length penalty of -0.5, using beam search with a size of 1. We utilize vLLM [37] as the backend to accelerate the generation process.
@@ 6.1 Main Results
Diagram Generation. Table 2 shows results on SVG-Diagrams, and Figures 4 and 14 provide visual examples. Results highlight that StarVector is the only method capable of performing diagram generation, as it uniquely applies the required primitives like rectangles, arrows, and text, whereas other methods produce blobs and curves that attempt to replicate structure and color. Metrics such as DinoScore, LPIPS, capture this advantage, while MSE and SSIM remain poorly aligned. Human evaluations further confirmed the preference for StarVector’s outputs.
@@ 7 Conclusions
Limitations and Future Work. StarVector is constrained by its 16k token context, which is inadequate for complex SVGs. It relies primarily on code prediction with minimal visual feedback. Additionally, generation speed is limited by the LLM. Future work may investigate integrating pixel-based signals to overcome these challenges.
Acknowledgements. We sincerely thank Ghazwa Darwiche, Christian Hudon, Fanny Rancourt, and Marie-Ève Marchand for their invaluable administrative and technical support. This work was supported by the Natural Sciences and Engineering Research Council of Canada and Mitacs. Chris Pal acknowledges the Canada CIFAR AI Chair. We also appreciate the human verifiers for their contributions to assessment and data verification.
@@ 8.3 Creating SVG-Diagrams
Figure 6: SVG Diagrams examples. These are ground truth SVG examples from the test set. They are presented as SVG, showing the challenge of understanding intricate structures and small texts, with images of variate aspect ratios.
@@ 10.0.1 Image-to-SVG Results
Comparing parameter count of models
. The number of learnable parameters in deep learning based models often correlates with performance. Pre-LLM models like Im2Vec and DeepSVG use significantly fewer parameters (up to 5M) compared to StarVector and GPT-based models, which operate in the billions. While pre-LLM models can produce accurate results, they lack the generalization ability of LLM-based approaches. Comparing StarVector with GPT models reveals that high-fidelity SVG generation is achievable with just 1B parameters, whereas GPT-4V lacks specific training for this task. Future models will likely incorporate SVG data in training, but current results already demonstrate that LLM-based approaches offer superior generalization and scalability for SVG generation, at the cost of utilizing more parameters.
Context Length Limitation.
The model’s architecture imposes a clear limitation on context length, which significantly impacts training and testing data pipelines, as well as the skills the model can learn. Our experiments show that the model scales effectively with increasing context lengths, from 8k to 16k, indicating that this limitation must be addressed with techniques for handling longer contexts—an area LLMs are expected to improve. For fair comparisons, we restricted our benchmark tests to a context length of 8k and evaluated all baselines within this setting. However, the benchmarks also provide versions with longer context lengths to assess future models, as increased length generally correlates with more complex SVGs. We did not observe substantial differences in scores between the 4k and 10k token settings, primarily because the data in our benchmark can typically be represented using an average of 3k tokens.
Limitations on Complex SVG Structures.
StarVector encounters challenges with complex SVG structures, intricate shapes, and detailed illustrations primarily due to limitations in its architecture. Currently, the model’s image encoder handles images by simply padding and resizing them to fixed dimensions of 224 or 384 pixels. This approach may not adequately capture the nuances of complex diagrams. A potential improvement would be to implement a dynamic image processing system akin to those found in newer Vision-Language Models (VLMs), which could enhance the model’s ability to interpret and generate intricate SVGs more effectively. Additionally, improving data cleaning processes is crucial, as the model sometimes produces hallucinated information due to noise in the input data, such as URLs or base64-encoded images. Addressing these issues through architectural enhancements and more robust data preprocessing could significantly improve StarVector’s performance on complex SVG tasks.
Generalization to Non-Standard SVGs.
StarVector’s ability to generalize to non-standard SVGs—those not represented in its main training distribution—poses a significant challenge. While the model performs well on common styles and primitives encountered during training, it struggles with more unique or unconventional SVGs. This is primarily due to the model’s training data, which tends to focus on widely used shapes and designs. As a result, StarVector may exhibit a bias towards these common styles, leading to suboptimal performance when faced with SVGs that feature unusual structures or less frequent elements.
To assess StarVector’s generalization capabilities, we evaluated its performance on various datasets that include non-standard SVGs. The results indicate that while the model can produce reasonable outputs for some non-standard examples, it often falls short in accurately capturing the intricacies of less familiar styles. This limitation suggests that the model’s training set lacks sufficient diversity to encompass the full range of potential SVG designs.
To address these concerns, future work should focus on expanding the training dataset to include a wider variety of SVG styles and structures. Incorporating data from niche applications and artistic domains could enhance the model’s ability to generate SVGs across a broader spectrum of design elements. Additionally, techniques such as domain adaptation and transfer learning could be explored to improve generalization to non-standard SVGs, allowing StarVector to adapt more effectively to unfamiliar inputs.